A quality control inspector found that 1.2% of 2500 telephones inspected were defective. How many telephones inspected were not defective?
Answer by richwmiller(17219)

multiply the percent by the number of phones.
1.2%*2500
convert 1.2% to a decimal or fraction
1.2%=1.2/100
1.2/100=.12/10=.012
1 percent (1/100) is .01 (one one hundredth)
so now we have .012*2500=30
It actually would have been easier to leave it as a fraction
1.2/100*2500=1.2*25=(1 1/5)*25=6/5*25=6*5=30